Basic scenario analysis - Prime Paints is in the process of evaluating two mutually exclusive additions to its processing capacity. The firm's financial analysts have developed pessimistic, most likely, and optimistic estimates of the annual cash inflows associated with each project. These estimates are shown in the following table:

Project A Project B
Initial Investment (CF0) $12,500 $12,500
Outcome Annual cash inflows (CF)
Pessimistic $900 $1,580
Most Likely $2,650 $1,650
Optimistic $2,490 $1,760

a. determine the range of annual cash inflows for each of the two projects.

b. assume that the firm's costs of capital is 9.2% and that both projects have 15-year lives. Construct a table showing the NPVs for each project for each of the possible outcomes. Include the range of NPVs for each project.

c. do parts (a) and (b) provide consistent views of the two projects? Explain

d. which project do you recommend? Why?
Answer #1

a. Range of annual cash inflows for each of the two projects is the difference between the highest annual CF and the lowest annual CF

  • Range for A = 2650-900 = 1750
  • Range for B = 1760-1580= 180

b. NPVs are calculated below:

Project A:

Project B

c. Both projects are consistent with the the answers in parts a and b because, as part a suggests, the range of CFs is very high in project A and so the NPV also varies a lot while the range is small in B and so NPV also doesn't vary much.

d. Recommendation can be made based on probability of each possibility. Suppose if each is equally likely then we can equally weight the NPVs

Project A

Project B

So as project A has a higher probability weighted NPV, it should be selected.